# tableTAB. 0

### v2

@@ VII.1 Dielectric Response of the Layered Plasmon System
$1-\frac{2\pi Ne^{2}q_{||}/m}{\omega\left(\omega+i\Gamma\right)-s^{2}q_{||}^{2}}\frac{\sinh(q_{||}d)}{\cosh(q_{||}d)-\cos(q_{\perp}d)}.$ (13)
@@ VII.3 EELS measurements
The momentum-resolved loss function was recorded at a FEI Titan3 TEM equipped with a Wien monochromator and a Gatan Tridiem imaging filter (GIF) at 80 kV acceleration voltage in a serial way (i.e., one EEL spectrum per fixed momentum). The energy resolution is around 120 meV (FWHM of zero loss peak). Using a camera length (i.e., effective distance between sample and detector) of 1.15 m. The GIF entrance aperture was used to select the different momentum values and covers a momentum range of 0.04 Å$\mbox{\AA}^{-1}$ (0.13  mrad semi collection angle). The acquisition times are presented in Table 1.
@@
VII.4 Fermi velocities and susceptibility
For the calculation of the Fermi velocity along the [100] direction and the bare particle susceptibility $\chi_{0}(q_{||},0,\omega)$ we used a TB band structure [47] based on a LDA band structure. We use a multi-band version of Eq. 2 and taking only intra-band excitations with the same matrix element into account, thus neglecting inter-band excitations which may be caused by a strong spin-orbit coupling, which leads to a k-dependence of the orbital character of the bands [49]. Mainly the $\beta$ and the $\gamma$ bands contribute to $v_{\text{F}}^{2}\right>_{100}$ and $\chi_{0}$.
@@ X Author contributions
J.F., M.K., A.L., and B.B. conceived the experiment, J.S. and D.W. performed the EELS experiment. J.S. and J.F. analyzed the data. F.J. and N.K. prepared and characterized the samples. J.F. calculated the susceptibility. J.F., J.S, and A.L. wrote the manuscript with input from all authors.

### v3

@@ VII.1 Dielectric Response of the Layered Plasmon System
$1-\frac{2\pi Ne^{2}q_{||}/m}{\omega\left(\omega+i\Gamma\right)-s^{2}q_{||}^{2}}\frac{\sinh(q_{||}d)}{\cosh(q_{||}d)-\cos(q_{\perp}d)}.$
@@ VII.3 EELS measurements
The momentum-resolved loss function was recorded at a FEI Titan3 TEM equipped with a Wien monochromator and a Gatan Tridiem imaging filter (GIF) at 80 kV acceleration voltage in a serial way (i.e., one EEL spectrum per fixed momentum). The energy resolution is around 120 meV (FWHM of zero loss peak). At a camera length (i.e., effective distance between sample and detector) of 1.15 m, the GIF entrance aperture was used to select the different momentum values and covers a momentum range of 0.04 Å$\mbox{\AA}^{-1}$ (0.13  mrad semi collection angle). The acquisition times are presented in Table 1.
@@
VII.4 Calculation of the average Fermi velocity, susceptibility, and loss function
For the calculation of the Fermi velocity along the [100] direction, the bare particle susceptibility $\chi_{0}(q_{||},\omega)$ and the loss function we used a TB band structure [47] based on an LDA band calculation. $\chi_{0}(q_{||},\omega)$ is calculated from a multi-band version of Eq. 2, taking only intra-band excitations with the same matrix element into account, thus neglecting inter-band excitations which may be caused by a strong spin-orbit coupling, which leads to a $k$-dependence of the orbital character of the bands [58]. Mainly the XY and the YZ bands contribute to $v_{\text{F}}^{2}\right>_{100}$ and $\chi_{0}$. The dielectric function and the loss function were calculated from Eqs. 1, 2, and 4. $\epsilon_{\text{b}}$ and the matrix element were fixed by using the plasmon energy from optical spectroscopy. For the calculation we used a half width of $\Gamma=$ 0.1 eV.
@@ X Author contributions
J.F., M.K., A.L., and B.B. conceived the experiment, J.S. and D.W. performed the EELS experiment. J.S. and J.F. analysed the data. F.J. and N.K. prepared and characterised the samples. J.F. calculated the susceptibility and the loss function. J.F., J.S, and A.L. wrote the manuscript with input from all authors.